What is the inverse of the function y=2ln(x-5)
My answer choices are:
A) y= (x/ln2)+5
B) x= (y/ln2)+5
C) y= 2e^(x-5)
D) y= e^(2x)+5
E) y= e^(0.5x)+5
I think it's A but I'm not sure. Please help!

Answer by stanbon(57387)

What is the inverse of the function y=2ln(x-5)
My answer choices are:
---
To find the inverse, interchange x and y to get:
x = 2ln(y-5)
Solve for "y":
ln(y-5) = x/2
y-5 = e^(x/2)
y = e^(x/2)+5
y = e^(0.5x) + 5
---
The answer is E.
Cheers,
Stan H.
-------------------




A) y= (x/ln2)+5
B) x= (y/ln2)+5
C) y= 2e^(x-5)
D) y= e^(2x)+5
E) y= e^(0.5x)+5